# A NICER View of PSR J1231--1411: A Complex Case

### v1

@@ 4.2 PDT-U
We present next the inference results for PSR J1231--1411 using the PDT-U model. In Figure 5 we show the radius, mass, compactness, and $\cos i$ posteriors for three radius priors: (1) using a lower limit of $8$ km (with the upper limit being $16$ km as usual); (2) limiting the radius between $10$ km and $14$ km, and (3) using an informative radius prior based on the results of Raaijmakers et al. (2021) (the R21 case from Section 3.3.2). The last was performed using both hydrogen (H) and helium (He) atmospheres (He R21 with 10,000 live points instead of the 20,000 used otherwise, but the H R21 results were checked and found to be close to identical with both settings). In all cases we report the $68.3\,\%$ credible intervals around the median. For the $R_{\mathrm{eq}}\in[8,16]$ km case we obtain $R_{\mathrm{eq}}=14.74_{-0.69}^{+0.62}$ km and $M=1.07_{-0.05}^{+0.06}$ $M_{\odot}$. For the $R_{\mathrm{eq}}\in[10,14]$ km case we obtain $R_{\mathrm{eq}}=13.52_{-0.50}^{+0.32}$ km and $M=1.06_{-0.04}^{+0.06}$ $M_{\odot}$. For the H R21 case we obtain $R_{\mathrm{eq}}=12.60_{-0.33}^{+0.31}$ km and $M=1.04_{-0.03}^{+0.05}$ $M_{\odot}$. Finally, for the He R21 case we obtain $R_{\mathrm{eq}}=12.48_{-0.34}^{+0.35}$ km and $M=1.09_{-0.06}^{+0.13}$ $M_{\odot}$.
@@
As mentioned also in Section 3.5, only the R21 and $R_{\mathrm{eq}}\in[10,14]$ km results are likely converged in terms of improving sampler settings. Both of them infer practically the same geometry solution (see the remaining posteriors in Appendix A) and yield almost equally good fit quality and evidence (see Table 1). On the contrary, the maximum likelihood and evidence values are significantly worse for the $R_{\mathrm{eq}}\in[8,16]$ km prior run: by $7$ and $17$, respectively (both in $\ln$-units) when comparing to the corresponding R21 run. This indicates that the wider radius prior run did not explore the parameter space sufficiently, and thus inferred a completely different (wrong) geometry solution. This was also the case in our early stage 4,000 live point run using the standard (broadest) radius prior for PDT-U (with some older settings). In that case both $5$ km and $15$ km radius modes were visibly detected in the posteriors  [^12] The former was better in the likelihood but the latter in the evidence., but both of them are significantly worse than the R21 and $R_{\mathrm{eq}}\in[10,14]$ km results in terms of likelihood and evidence. From Table 1 we also see that the evidence and maximum likelihood are highest for the He R21 run: by $8$ and $3$ in $\ln$-units, respectively, compared to the H R21 run. All of the credible intervals for the two atmosphere options are, however, very consistent with each other (with He giving often broader intervals).
@@
The performance of the PDT-U model was found to be good, both in terms of two-dimensional NICER residuals (see Figure 6) and bolometric $\chi^{2}$ value of the maximum likelihood parameter vector (see the right panel of Figure 4). The $\chi^{2}$ value was around $18$ for the He R21 run, $19-20$ for the H R21 and $R_{\mathrm{eq}}\in[10,14]$ km runs, and $21$ for the non-converged H $R_{\mathrm{eq}}\in[8,16]$ km run (all again with $32$ phase bins). We also checked if the inferred NICER background (based mostly on simultaneous XMM-Newton fits) is consistent with the predicted SCORPEON background in Figure 7. We see that it generally is, except at the lowest energies (close to $0.3$ keV) where the inferred background is somewhat below, at $0.4-0.6$ keV where it is somewhat above, and at $1.0-1.4$ keV where it is slightly below the SCORPEON range. Inferring a slightly higher background is not surprising, since the SCORPEON estimate does not include any unpulsed and unmodeled emission from the NS nor any emission from nearby sources (see the XMM-Newton image in Figure 8 of Bogdanov et al. 2019a). In addition, as noted in Section 2.1, the SCORPEON background estimate was found to depend on the chosen source spectral model (especially at the lowest energies), and none of the tried XSPEC models matches exactly the model used in X-PSI and its inferred parameters. Therefore, we consider the agreement reasonable. We also note that in our early NICER-only runs (without XMM-Newton) the inferred background was found to be substantially too large (covering all the unpulsed signal).
@@
The contribution from the two different hot regions to the total pulse can be viewed from Figure 8, both for the incident signal and for that registered by NICER (without background). We see that the pulse from the primary region has a smoother shape than that from the secondary region, which remains mostly flat between the pulse peaks. The corresponding inferred hot region geometry is shown in Figure 9 for the maximum likelihood sample of the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km run. The primary region is located on the northern hemisphere and consists of a rather small hot component surrounded by a very large cold component. The secondary region is located close to the southern pole and consists of almost equally hot but barely connected components with similar sizes. We note that the maximum likelihood geometry is very similar for the PDT-U H $R_{\mathrm{eq}}$ R21 and PDT-U He $R_{\mathrm{eq}}$ R21 runs, except that the small hot component of the primary region is located at the equator in the latter case (which is still consistent with the H cases considering the width of the credible intervals).
@@
The inferred values for the remaining parameters, e.g., for $D$ and $N_{\mathrm{H}}$, are found to be well consistent with expectations (see Section 5.1). For all the models we find the distance to be around $D=0.65\pm 0.05$ kpc. Inferred column densities show slightly more scatter, but all of the converged PDT-U runs provide roughly $N_{\mathrm{H}}=1.8\pm 0.5\times 10^{20}\mathrm{cm}^{-2}$.
@@ 5.1 Implications
One of the more robust findings of this work is that the preferred hot region model is PDT-U rather than ST-U. This means that assuming two circular uniformly emitting hot regions is not sufficient and may lead to wrong estimates of the NS parameters. Assuming a NS radius consistent with the previous EOS analysis and using the PDT-U model, the inferred hot region geometry is also constrained to be rather non-antipodal (see Figure 9), which implies that purely centered dipolar magnetic field is unlikely for PSR J1231--1411 (as for other NICER RMPs analyzed so far). The large cold emission component surrounding the primary hot spot may also be associated with emission from the rest of the star’s surface, hinting that using a ST+PDT model (see Vinciguerra et al. 2024 for a definition) and allowing the rest of the star surface to emit might lead to similar results. However, this is expected to speed-up the computation only mildly, and was thus not tried here.
@@
We can also note a few interesting details about the inferred mass and radius. We see from Figure 5 that all of the inferred masses are quite small, and close to our $1$ $M_{\odot}$ lower limit (that limit is motivated by supernova theory, see Müller et al. 2024, and by the description of the early evolution of a NS, see Strobel et al. 1999). These mass constraints are notably tighter than the mass prior which we used from radio timing measurements, but they are still in the region where the prior is highest (see Figure 3). The radius constraints, on the other hand, do not provide as much information gain over the prior, especially in the R21 case. There the Kullback–Leibler (KL) divergence [^13] Representing prior-to-posterior information gain. is $0.55$, implying a non-zero information given by the data. In the H $R_{\mathrm{eq}}\in[10,14]$ km case, the KL-divergence of radius is notably better ($1.40$) but still worse than that of mass ($2.55$).
@@
Additionally, our results offer some insights into the NS atmosphere composition by finding better evidences and likelihoods for helium rather than hydrogen when using the PDT-U model and R21 radius prior (see Section 4.2, the same was however also detected in preliminary non-restricted ST-U analyses). However, the difference is not large enough to be conclusive, considering that hydrogen composition is a priori more probable based on rapid sinking of heavier elements via diffusive gravitational separation (Alcock & Illarionov, 1980; Hameury et al., 1983; Brown et al., 2002; Zavlin & Pavlov, 2002). A helium atmosphere could occur if all the hydrogen was converted to helium via diffusive nuclear burning (Wijngaarden et al., 2019), or if the star had only accreted from a completely hydrogen-depleted companion star (see Bogdanov et al., 2021; Salmi et al., 2023, and the references therein). Nevertheless, there is evidence that the companion of PSR J1231--1411 is a white dwarf with a hydrogen envelope (Bassa et al., 2016).
@@
The $N_{\mathrm{H}}$ column density can be estimated via independent means and compared to the inferred values. Neutral hydrogen maps predict values $\sim 3.4\times 10^{20}\,\mathrm{cm}^{-2}$ (HI4PI Collaboration et al., 2016), but this should be considered an approximate upper limit as this represent the column density integrated to the edge of the Galaxy in that direction [^14] For this, we used the HEASARC $N_{\mathrm{H}}$ tool at https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3nh/w3nh.pl.. Other estimates of $N_{\mathrm{H}}$, at the distance of the pulsar, can be obtained from 3-dimensional maps of infrared or optical extinction $A_{V}$ or reddening $E\left(B-V\right)$. At the position and distance of PSR J1231--1411, $\sim 0.6$ kpc, we find $\sim 3.5\times 10^{20}\,\mathrm{cm}^{-2}$ based on the 3D dust maps of Green et al. (2019)  [^15] Query tool is available at http://argonaut.skymaps.info/. (this 3D map shows little dependence of $E\left(B-V\right)$ on the distance beyond 0.2 kpc). Alternatively, with the 3D dust maps of Lallement et al. (2022); Vergely et al. (2022)  [^16] Data are available on Vizier at (https://vizier.cds.unistra.fr) or via an online tool at https://explore-platform.eu/sdas/about/gtomo., we find $\sim 3.4\times 10^{20}\,\mathrm{cm}^{-2}$. Both results make use of the relation of Foight et al. (2016) between $N_{\mathrm{H}}$ and $A_{V}$. Another estimate can exploit the relation between the dispersion measure (DM, number density of free electrons in the line of sight) and $N_{\mathrm{H}}$ (He et al., 2013). For PSR J1231--1411, ${\rm DM}=8.09\,\mathrm{pc}\,\mathrm{cm}^{-3}$, which gives $\sim 2.4\times 10^{20}\,\mathrm{cm}^{-2}$.
@@
The $N_{\mathrm{H}}$ values inferred in this work are globally consistent with these independent measurements. It is important to keep in mind that the measurements deduced above should be considered carefully since the employed relations may be uncertain, especially at low-absorption values because the sampling of objects is much smaller for low values of $N_{\mathrm{H}}$. In other words, these independent measurements mostly bring support that the chosen prior on $N_{\mathrm{H}}$ is broad enough and does not bias the inferred parameters.
@@
Our inferred distance is also found to be well consistent with the preliminary parallax distance measurement (Cromartie et al., 2024), both being around or above $600$ pc. We note though that our distance upper limit ($700$ pc) might be too restrictive, since our posteriors peak only slightly below it and the parallax distance uncertainties can also allow higher distances, especially when accounting for the Lutz-Kelker bias (Lutz & Kelker, 1973; Verbiest et al., 2012). Using a more sophisticated distance prior is, however, left to future work.
@@ 5.2 Comparison to Other Pulsars
Compared to the other NICER RMPs analyzed so far, constraining the NS parameters for PSR J1231--1411 appears more challenging. There are several possible reasons for this. First of all, the available priors on mass and observer inclination are much less restrictive for PSR J1231--1411 than for PSR J0740$+$6620 and PSR J0437--4715. For the isolated pulsar PSR J0030$+$0451 there is even less prior information about mass and inclination, but the distance is known much more precisely based on radio timing parallax measurements (Arzoumanian et al., 2018). These factors, together with the relatively large number of free hot region geometry parameters needed for the PDT-U model, likely make the sampling problem more difficult. In the previous X-PSI analyses, the simple ST-U model was found to be adequate for PSR J0740$+$6620 (Riley et al., 2021) and the CST+PDT model, with three parameters less than for PDT-U, for PSR J0437--4715 (Choudhury et al., 2024a). In Vinciguerra et al. (2024) PDT-U was one of the models applied to PSR J0030$+$0451, however its convergence was not investigated.
PSR J1231--1411 differs from the other pulsars also in terms of the pulse shape. As seen from Figures 1 and 8, there is only a weak interpulse appearing in the bolometric pulse profile. This feature is especially weak in the two-dimensional data leading possibly to multitude of solutions that can more or less fit it, but that are too distinct in the parameter space to be explored thoroughly if one does not shrink the prior space. For PSR J0030$+$0451 and PSR J0740$+$6620 the secondary pulse is much stronger, and for PSR J0437--4715 there is no peaky interpulse (see Bogdanov et al. 2019a and Wolff et al. 2021). However, it could also be that the source properties of PSR J1231--1411 simply differ from the others in such a way that they are more difficult to recover.
It is nevertheless encouraging that despite the uncertainty, we are able to obtain solutions for this source that fit the data well and are also consistent with the results from other NICER RMPs and multi-messenger analyses. There are clearly various steps that might improve the situation: tighter priors, additional computational resources, or a longer data set with more photons. However before devoting additional resources (in the form of observing or computational time) to this source, we would recommend a programme of simulations using synthetic data to determine whether this is likely to lead to improvements, or whether there are fundamental issues with this type of pulse profile that are likely to prevent efforts to obtain tighter and more robust constraints.

### v2

@@ 4.2 PDT-U
We present next the inference results for PSR J1231--1411 using the PDT-U model. In Figure 5 we show the radius, mass, compactness, and $\cos i$ posteriors for three radius priors: (1) using a lower limit of $8$ km (with the upper limit being $16$ km as usual); (2) limiting the radius between $10$ km and $14$ km, and (3) using an informative radius prior based on the results of Raaijmakers et al. (2021, the R21 case from Section 3.3.2). The last was performed using both hydrogen (H) and helium (He) atmospheres (He R21 with 10,000 live points instead of the 20,000 used otherwise, but the H R21 results were checked and found to be close to identical with both settings). In all cases, we report the $68.3\,\%$ credible intervals around the median. For the $R_{\mathrm{eq}}\in[8,16]$ km case, we obtain $R_{\mathrm{eq}}=14.74_{-0.69}^{+0.62}$ km and $M=1.07_{-0.05}^{+0.06}$ $M_{\odot}$. For the $R_{\mathrm{eq}}\in[10,14]$ km case, we obtain $R_{\mathrm{eq}}=13.52_{-0.50}^{+0.32}$ km and $M=1.06_{-0.04}^{+0.06}$ $M_{\odot}$. For the H R21 case, we obtain $R_{\mathrm{eq}}=12.60_{-0.33}^{+0.31}$ km and $M=1.04_{-0.03}^{+0.05}$ $M_{\odot}$. Finally, for the He R21 case, we obtain $R_{\mathrm{eq}}=12.48_{-0.34}^{+0.35}$ km and $M=1.09_{-0.06}^{+0.13}$ $M_{\odot}$.
@@
As mentioned also in Section 3.5, only the R21 and $R_{\mathrm{eq}}\in[10,14]$ km results are likely converged in terms of improving sampler settings. Both of them infer practically the same geometry solution (see the remaining posteriors in Appendix A) and yield almost equally good fit quality and evidence (see Table 1). On the contrary, the maximum-likelihood and evidence values are significantly worse for the $R_{\mathrm{eq}}\in[8,16]$ km prior run: by $7$ and $17$, respectively (both in $\ln$-units) when comparing to the corresponding R21 run. This indicates that the wider radius prior run did not explore the parameter space sufficiently, and thus inferred a completely different (wrong) geometry solution. This was also the case in our early stage 4000 live point run using the standard (broadest) radius prior for PDT-U (with some older settings). In that case, both $5$ km and $15$ km radius modes were visibly detected in the posteriors, [^25] The former was better in the likelihood but the latter in the evidence. but both of them are significantly worse than the R21 and $R_{\mathrm{eq}}\in[10,14]$ km results in terms of likelihood and evidence. From Table 1, we also see that the evidence and maximum likelihood are highest for the He R21 run: by $8$ and $3$ in $\ln$-units, respectively, compared to the H R21 run. All of the credible intervals for the two atmosphere options are, however, very consistent with each other (with He giving often broader intervals).
@@
The performance of the PDT-U model was found to be good, both in terms of the 2D NICER residuals (see Figure 6) and the bolometric $\chi^{2}$ value of the maximum-likelihood parameter vector (see the right-hand panel of Figure 4). The $\chi^{2}$ value was around $18$ for the He R21 run, $19-20$ for the H R21 and $R_{\mathrm{eq}}\in[10,14]$ km runs, and $21$ for the nonconverged H $R_{\mathrm{eq}}\in[8,16]$ km run (all again with $32$ phase bins). We also checked if the inferred NICER background (based mostly on simultaneous XMM-Newton fits) is consistent with the predicted SCORPEON background in Figure 7. We see that it generally is, except at the lowest energies (close to $0.3$ keV) where the inferred background is somewhat below, at $0.4-0.6$ keV where it is somewhat above, and at $1.0-1.4$ keV where it is slightly below the SCORPEON range. Inferring a slightly higher background is not surprising because the SCORPEON estimate does not include any unpulsed and unmodeled emission from the NS nor any emission from nearby sources (see the XMM-Newton image in Figure 8 of Bogdanov et al. 2019a). In addition, as noted in Section 2.1, the SCORPEON background estimate was found to depend on the chosen source spectral model (especially at the lowest energies), and none of the tried XSPEC models matches exactly the model used in X-PSI and its inferred parameters. Therefore, we consider the agreement reasonable. We also note that in our early NICER-only runs (without XMM-Newton) the inferred background was found to be substantially too large (covering all the unpulsed signal).
@@
The contribution from the two different hot regions to the total pulse can be viewed from Figure 8, both for the incident signal and for that registered by NICER (without background). We see that the pulse from the primary region has a smoother shape than that from the secondary region, which remains mostly flat between the pulse peaks. The corresponding inferred hot region geometry is shown in Figure 9 for the maximum-likelihood sample of the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km run. The primary region is located on the northern hemisphere and consists of a rather small hot component surrounded by a very large cold component. The secondary region is located close to the southern pole and consists of almost equally hot but barely connected components with similar sizes. We note that the maximum-likelihood geometry is very similar for the PDT-U H $R_{\mathrm{eq}}$ R21 and PDT-U He $R_{\mathrm{eq}}$ R21 runs, except that the small hot component of the primary region is located at the equator in the latter case (which is still consistent with the H cases considering the width of the credible intervals).
@@
The inferred values for the remaining parameters, e.g., for $D$ and $N_{\mathrm{H}}$, are found to be well consistent with expectations (see Section 5.1). For all the models, we find the distance to be around $D=0.65\pm 0.05$ kpc. Inferred column densities show slightly more scatter, but all of the converged PDT-U runs provide roughly $N_{\mathrm{H}}=1.8\pm 0.5\times 10^{20}\mathrm{cm}^{-2}$.
@@ 5.1 Implications
One of the more robust findings of this work is that the preferred hot region model is PDT-U rather than ST-U. This means that assuming two circular uniformly emitting hot regions is not sufficient and may lead to wrong estimates of the NS parameters. Assuming a NS radius consistent with the previous EOS analysis and using the PDT-U model, the inferred hot region geometry is also constrained to be rather nonantipodal (see Figure 9), which implies that purely centered dipolar magnetic field is unlikely for PSR J1231--1411 (as for other NICER RMPs analyzed so far). The large cold emission component surrounding the primary hot spot may also be associated with emission from the rest of the star’s surface, hinting that using a ST+PDT model (see Vinciguerra et al. 2024 for a definition) and allowing the rest of the star surface to emit might lead to similar results. However, this is expected to speed-up the computation only mildly, and was thus not tried here.
@@
We can also note a few interesting details about the inferred mass and radius. We see from Figure 5 that all of the inferred masses are quite small and close to our $1$ $M_{\odot}$ lower limit (this limit is motivated by supernova theory, see Müller et al. 2024, and by the description of the early evolution of a NS, see Strobel et al. 1999). These mass constraints are notably tighter than the mass prior which we used from radio timing measurements but they are still in the region where the prior is highest (see Figure 3). The radius constraints, on the other hand, do not provide as much information gain over the prior, especially in the R21 case. There the Kullback–Leibler (KL) divergence [^26] Representing prior-to-posterior information gain. is $0.55$, implying a nonzero information given by the data. In the H $R_{\mathrm{eq}}\in[10,14]$ km case, the KL-divergence of radius is notably better ($1.40$) but still worse than that of mass ($2.55$).
@@
Additionally, our results offer some insights into the NS atmosphere composition by finding better evidence and likelihoods for helium rather than hydrogen when using the PDT-U model and R21 radius prior (see Section 4.2, the same was, however, also detected in preliminary nonrestricted ST-U analyses). However, the difference is not large enough to be conclusive considering that hydrogen composition is a priori more probable based on the rapid sinking of heavier elements via diffusive gravitational separation (Alcock & Illarionov, 1980; Hameury et al., 1983; Brown et al., 2002; Zavlin & Pavlov, 2002). A helium atmosphere could occur if all the hydrogen was converted to helium via diffusive nuclear burning (Wijngaarden et al., 2019), or if the star had only accreted from a completely hydrogen-depleted companion star (see Bogdanov et al., 2021; Salmi et al., 2023, and the references therein). Nevertheless, there is evidence that the companion of PSR J1231--1411 is a white dwarf with a hydrogen envelope (Bassa et al., 2016).
@@
The $N_{\mathrm{H}}$ column density can be estimated via independent means and compared to the inferred values. Neutral hydrogen maps predict values $\sim 3.4\times 10^{20}\,\mathrm{cm}^{-2}$ (HI4PI Collaboration et al., 2016), but this should be considered an approximate upper limit because this represents the column density integrated to the edge of the Galaxy in that direction. [^27] For this, we used the HEASARC $N_{\mathrm{H}}$ tool at https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3nh/w3nh.pl. Other estimates of $N_{\mathrm{H}}$, at the distance of the pulsar, can be obtained from 3D maps of infrared or optical extinction $A_{V}$ or reddening $E\left(B-V\right)$. At the position and distance of PSR J1231--1411, $\sim 0.6$ kpc, we find $\sim 3.5\times 10^{20}\,\mathrm{cm}^{-2}$ based on the 3D dust maps of Green et al. (2019,  [^28] Query tool is available at http://argonaut.skymaps.info/. this 3D map shows little dependence of $E\left(B-V\right)$ on the distance beyond 0.2 kpc). Alternatively, with the 3D dust maps of Lallement et al. (2022) and Vergely et al. (2022), [^29] Data are available on Vizier at (https://vizier.cds.unistra.fr) or via an online tool at https://explore-platform.eu/sdas/about/gtomo. we find $\sim 3.4\times 10^{20}\,\mathrm{cm}^{-2}$. Both results make use of the relation of Foight et al. (2016) between $N_{\mathrm{H}}$ and $A_{V}$. Another estimate can exploit the relation between the DM (number density of free electrons in the line of sight) and $N_{\mathrm{H}}$ (He et al., 2013). For PSR J1231--1411, ${\rm DM}=8.09\,\mathrm{pc}\,\mathrm{cm}^{-3}$, which gives $\sim 2.4\times 10^{20}\,\mathrm{cm}^{-2}$.
@@
The $N_{\mathrm{H}}$ values inferred in this work are globally consistent with these independent measurements. It is important to keep in mind that the measurements deduced above should be considered carefully because the employed relations may be uncertain, especially at low-absorption values because the sampling of objects is much smaller for low values of $N_{\mathrm{H}}$. In other words, these independent measurements mostly bring support that the chosen prior on $N_{\mathrm{H}}$ is broad enough and does not bias the inferred parameters.
@@
Our inferred distance is also found to be well consistent with the preliminary parallax distance measurement (Cromartie et al. 2024, in preparation), both being around or above $600$ pc. We note though that our distance upper limit ($700$ pc) might be too restrictive because our posteriors peak only slightly below it and the parallax distance uncertainties can also allow higher distances, especially when accounting for the Lutz-Kelker bias (Lutz & Kelker, 1973; Verbiest et al., 2012). Using a more sophisticated distance prior is, however, left to future work.
@@ 5.2 Comparison to Other Pulsars
Compared to the other NICER RMPs analyzed so far, constraining the NS parameters for PSR J1231--1411 appears more challenging. There are several possible reasons for this. First of all, the available priors on mass and observer inclination are much less restrictive for PSR J1231--1411 than for PSR J0740$+$6620 and PSR J0437--4715. For the isolated pulsar PSR J0030$+$0451, there is even less prior information about mass and inclination, but the distance is known much more precisely based on radio timing parallax measurements (Arzoumanian et al., 2018). These factors, together with the relatively large number of free hot region geometry parameters needed for the PDT-U model, likely make the sampling problem more difficult. In the previous X-PSI analyses, the simple ST-U model was found to be adequate for PSR J0740$+$6620 (Riley et al., 2021) and the CST+PDT model, with three parameters less than for PDT-U, for PSR J0437--4715 (Choudhury et al., 2024a). In Vinciguerra et al. (2024) PDT-U was one of the models applied to PSR J0030$+$0451, however its convergence was not investigated.
PSR J1231--1411 also differs from the other pulsars also in terms of the pulse shape. As seen from Figures 1 and 8, there is only a weak interpulse appearing in the bolometric pulse profile. This feature is especially weak in the 2D data leading possibly to a multitude of solutions that can more or less fit it but that are too distinct in the parameter space to be explored thoroughly if one does not shrink the prior space. For PSR J0030$+$0451 and PSR J0740$+$6620, the secondary pulse is much stronger, and for PSR J0437--4715 there is no peaky interpulse (see Bogdanov et al. 2019a and Wolff et al. 2021). However, it could also be that the source properties of PSR J1231--1411 simply differ from the others in such a way that they are more difficult to recover.
It is nevertheless encouraging that despite the uncertainty, we are able to obtain solutions for this source that fit the data well and are also consistent with the results from other NICER RMPs and multimessenger analyses. There are clearly various steps that might improve the situation: tighter priors, additional computational resources, or a longer data set with more photons. However, before devoting additional resources (in the form of observing or computational time) to this source, we would recommend a program of simulations using synthetic data to determine whether this is likely to lead to improvements or whether there are fundamental issues with this type of pulse profile that are likely to prevent efforts to obtain tighter and more robust constraints.